what is the cube root of 1215
Answer by CharlesG2(828)

what is the cube root of 1215
you could just use a calculator you know, or do you want it factored out?
1215 = 1 * 1215 = 1 * 3 * 405 = 1 * 3 * 3 * 135 = 1 * 3 * 3 * 3 * 45
= 1 * 3 * 3 * 3 * 3 * 15 = 1 * 3 * 3 * 3 * 3 * 3 * 5
= 1 * 3^3 * 3^2 * 5
= 3 * cube root of 45
you can use a calculator now if you need to go further